OpenStudy (anonymous):

f(x)=-15x+1 g(x)=-15x^2-20x-5 find (gof)(-3)

OpenStudy (anonymous):

woah that looks confuzing, okay: f(x)=15x+1 g(x)=-15x\[^{2}\]-20x-5 find (g o f)(3)

OpenStudy (anonymous):

*-3, sorry

hero (hero):

f(x)=-15x+1 g(x)=-15x^2-20x-5 f(-3) = -15(-3) + 1 = 45 + 1 = 46 (g o f)(-3) = g(f(-3)) = g(46) g(46) = -15(46)^2 - 20(46) - 5 = -31740 - 920 - 5 = -32665

hero (hero):

That should be the correct answer

OpenStudy (anonymous):

dude you're awesome. i remember that now. so you solve f(-3) and then put it into g ?

hero (hero):

The general formula is: (f o g)(x) = f(g(x)) But in this case we have: (g o f)(-x), so we write: (g o f)(-x) = g(f(-x)), then solve.

OpenStudy (anonymous):

sweethanks !!! could you help me with one more problem?

hero (hero):

Post it